# Large Deviations Principle for Isoperimetry and Its Equivalence to Nonlinear Log-Sobolev Inequalities

### v3

@@ 3.1. From Transport to Variant Isoperimetry
Case of $\epsilon\leq d(x,y)$: For $x^{\prime}$ such that $d(x,x^{\prime})\leq\epsilon$, we have
(3.9) $d^{p}(x^{\prime},y)$ $\geq(d(x,y)-d(x^{\prime},x))^{p}\geq(d(x,y)-\epsilon)^{p}\geq\frac{1}{(1+\epsilon)^{p-1}}d^{p}(x,y)-\epsilon,$
where the last inequality follows by Jensen’s inequality:
$\frac{1}{1+\epsilon}\cdot(d(x,y)-\epsilon)^{p}+\frac{\epsilon}{1+\epsilon}\cdot 1^{p}\geq\left(\frac{d(x,y)}{1+\epsilon}\right)^{p}.$
We set $\delta(\epsilon)=\max\{1-\frac{1}{(1+\epsilon)^{p-1}},\epsilon\}$. Then, the inequalities in (3.9) imply (3.8).
Case of $\epsilon>d(x,y)$: As shown above, the inequality in (3.8) holds for $\epsilon=d(x,y)$. Note that the right hand side of (3.8) is increasing in $d(x,y)$ and the LHS is always zero for the case $d(x,y)\leq\epsilon$. So, for all $d(x,y)\leq\epsilon$, (3.8) still holds. ∎
Based on the above notation, we introduce the following quantity $\psi$ that is particularly important in large deviations analysis for isoperimetry.
$\psi(\alpha,\tau):=\psi(\alpha,\tau|\nu_{X},\nu_{Y})$ $:=\sup_{\pi_{XW}\in P_{\mathrm{c}}(\mathcal{X}\times\{0,1\}):D(\pi_{X|W}\|\nu_{X}|\pi_{W})\leq\alpha}$
(3.10) $\qquad\inf_{\pi_{Y|W}:\mathsf{C}(\pi_{X|W},\pi_{Y|W}|\pi_{W})\leq\tau}D(\pi_{Y|W}\|\nu_{Y}|\pi_{W}),$
where $W$ is defined on $\{0,1\}$, and $P_{\mathrm{c}}(\mathcal{X}\times\{0,1\})$ denotes the set of probability measures concentrated on compact sets. The infimum above actually represents a nonlinear transport inequality. By the convexity of optimal transport cost and the relative entropy, one can easily see that $\tau\mapsto\psi(\alpha,\tau)$ is convex. Define
(3.11) $\psi_{\epsilon,\kappa}(\alpha,\tau)$ $:=\psi_{\epsilon,\kappa}(\alpha,\tau|\nu_{X},\nu_{Y})$
(3.12) $:=\sup_{\pi_{XW}\in P_{\mathrm{c}}(\mathcal{X}\times\{0,1\}):D(\pi_{X|W}\|\nu_{X}|\pi_{W})\leq\alpha}$
$\qquad\inf_{\pi_{Y|XW}:\mathbb{E}_{\pi}[c(X,Y)]\leq\tau,\,\mathbb{E}_{\pi}[c^{1+\epsilon}(X,Y)]\leq\kappa}D(\pi_{Y|W}\|\nu_{Y}|\pi_{W}).$
Without changing the optimal value, the range of $W$ in (3.11) can be changed to any larger but finite set. This point can be proven by following similar proof steps as [68, Theorem 20].
Based on $\psi$ and $\psi_{\epsilon,\kappa}$, the asymptotic expression of $E_{n}$ is characterized in our previous work.
Theorem 8 (Asymptotic Exponent [68, Theorem 12]).
The following hold.
  1. (1)
    Assume that $\nu_{X}$ is compactly supported and Assumption 3 holds. Then, for any $\alpha\geq 0,\tau>0$, it holds that
    (3.13) $\limsup_{n\to\infty}E_{n}(\alpha,\tau)\leq\lim_{\epsilon\downarrow 0,\kappa\uparrow\infty}\psi_{\epsilon,\kappa}(\alpha,\tau),$
    where the limit at the RHS can be replaced by $\inf_{\epsilon>0,\kappa>0}$.
  2. (2)
    Assume that $\mathcal{X}=\mathcal{Y}$ equipped with a metric $d$ is a Polish metric space, and the cost function is set to $c=d^{p}$ for $1\leq p<\infty$. Then, for any $(\alpha,\tau)$ in the interior of $\mathrm{dom}\psi:=\left\{(\alpha,\tau):\psi(\alpha,\tau)<\infty\right\}$, it holds that
    $\liminf_{n\to\infty}E_{n}(\alpha,\tau)\geq\psi(\alpha,\tau).$
Remark 7.
Statement 1 in this theorem follows from the original proof of Statement 1 in [68, Theorem 12] with two slight modifications. Firstly, we replace the power 2 in the equation chain in (56) of [68] by $q=1+\epsilon$, and then apply the Marcinkiewicz–Zygmund inequality to upper bound $\mathbb{E}_{\pi}\left[|c_{n}(x^{n},Y^{n})-n\mu|^{q}\right]$ in terms of $\mathbb{E}_{\pi}\left[c(X,Y)^{q}\right]$ (or $\kappa$). The Marcinkiewicz–Zygmund inequality here refers to that for $1<q<2$, $\mathbb{E}\left[|\sum_{i=1}^{n}X_{i}|^{q}\right]\leq B_{q}\mathbb{E}\left[(\sum_{i=1}^{n}|X_{i}|^{2})^{q/2}\right]\leq B_{q}\sum_{i=1}^{n}\mathbb{E}\left[|X_{i}|^{q}\right]$ for independent $X_{i}$’s such that $\mathbb{E}\left[X_{i}\right]=0$, where $B_{q}$ is a constant only depending on $q$. Secondly, Assumption 3 is modified in the present form in order to include Example 1 (with unbounded metrics). Statement 2 is also given in [68, Theorem 12].
Remark 8.
By checking the proof, Theorem 8 still holds if the constraint $\pi_{XW}\in P_{\mathrm{c}}(\mathcal{X}\times\{0,1\})$ is replaced by $\pi_{XW}\in P(\mathcal{X}\times\{0,1\})$.
The inherently typical subset lemma in [2] plays a key role in the proof of the theorem above, which enables us to tailor a set into a new one that is “regular” in some sense, so that we can analyze the asymptotic expression of the boundary size of the set.
Why is the asymptotic exponent determined by the relative entropy and the Wasserstein metric? The intuition is as follows. For a distribution $\nu$, denote $B_{\epsilon]}(\nu)$ as the closed ball of center $\nu$ and radius $\epsilon$ under the Lévy–Prokhorov metric. Let $\mathsf{L}_{n}:x\in\mathcal{X}^{n}\mapsto\mathsf{L}_{x}\in P(\mathcal{X})$ be the empirical measure map, where $\mathsf{L}_{x}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}}$ is the empirical measure of $x$, and $P(\mathcal{X})$ is the set of probability measures on $\mathcal{X}$. For a Polish space $\mathcal{X}$ and $\epsilon>0$, the empirically $\epsilon$-typical set of $\pi$ [47] is defined as
(3.14) $\mathcal{T}_{\epsilon}^{(n)}(\pi):=\mathsf{L}_{n}^{-1}(B_{\epsilon]}(\pi)).$
For an empirically $\epsilon$-typical set $A=\mathsf{L}_{n}^{-1}(B_{\epsilon}(\pi_{X}))$ (i.e., $A=\{\mathbf{x}\in\mathcal{X}^{n}:\mathsf{L}_{\mathbf{x}}\approx\pi_{X}\}$), roughly speaking, by Sanov’s theorem it holds that $\nu_{n}(A)\approx e^{-nD(\pi_{X}\|\nu)}$ for $\epsilon$ chosen sufficiently small, which is no smaller than $e^{-n\alpha}$ if $D(\pi_{X}\|\nu)\leq\alpha$. The enlargement of this set $A^{c_{n}\leq n\tau}$ can be written as
$A^{c_{n}\leq n\tau}$ $=\{\mathbf{y}:\frac{1}{n}c_{n}(\mathbf{x},\mathbf{y})\leq\tau,\exists\mathbf{x}\in A\}$
$=\{\mathbf{y}:\mathbb{E}_{\mathsf{L}_{\mathbf{x},\mathbf{y}}}[c(X,Y)]\leq\tau,\mathsf{L}_{\mathbf{x}}\approx\pi_{X},\exists\mathbf{x}\}$
(3.15) $=\{\mathbf{y}:\mathsf{C}(\mathsf{L}_{\mathbf{x}},\mathsf{L}_{\mathbf{y}})\leq\tau,\mathsf{L}_{\mathbf{x}}\approx\pi_{X}\}$
$\approx\mathsf{L}^{-1}\{\pi_{Y}:\mathsf{C}(\pi_{X},\pi_{Y})\leq\tau\},$
where (3.15) is due to Birkhoff’s theorem, $\min_{\sigma}\mathbb{E}_{\mathsf{L}_{\sigma(\mathbf{x}),\mathbf{y}}}[c(X,Y)]=\mathsf{C}(\mathsf{L}_{\mathbf{x}},\mathsf{L}_{\mathbf{y}})$, with $\sigma$ denoting a permutation operator [64, Page 5]. By Sanov’s theorem again,
$\nu^{\otimes n}(A^{c_{n}\leq n\tau})\approx\exp\{-n\inf_{\pi_{Y}:\mathsf{C}(\pi_{X},\pi_{Y})\leq\tau}D(\pi_{Y}\|\nu)\}.$
Optimizing over $\pi_{X}$ such that $D(\pi_{X}\|\nu)\leq\alpha$, we obtain the isoperimetric exponent for $A$,
$\sup_{\pi_{X}:D(\pi_{X}\|\nu)\leq\alpha}\inf_{\pi_{Y}:\mathsf{C}(\pi_{X},\pi_{Y})\leq\tau}D(\pi_{Y}\|\nu).$
By product construction of such $A$’s (forming a conditional empirically typical set), we obtain the exponent $\psi(\alpha,\tau)$. That is, conditional empirically $\epsilon$-typical sets are optimal in attaining the optimal exponent.
Statement 1 in Theorem 8 holds for Example 5 since Assumption 3 is satisfied in this case.
Theorem 9 (Asymptotic Exponent for Metric-Valued Cost and Compactly Supported $\nu_{X}$).
Assume that $\mathcal{X}=\mathcal{Y}$ equipped with a metric $d$ is a Polish metric space, and the cost function is set to $c=d^{p}$ for $1\leq p<\infty$. Assume that $\nu_{X}$ is compactly supported and $\mathbb{E}_{\nu_{Y}}[d(Y,x_{0})^{q}]<\infty$ for some $q>p$ and $x_{0}\in\mathcal{X}$. Then, for any $(\alpha,\tau)$ in the interior of $\mathrm{dom}\psi$,
(3.16) $\lim_{n\to\infty}E_{n}(\alpha,\tau)=\psi(\alpha,\tau),$
where
$\psi(\alpha,\tau)=\psi(\alpha,\tau|\nu_{X},\nu_{Y})$ $=\sup_{\pi_{XW}\in P_{\mathrm{c}}(\mathcal{X}\times\{0,1\}):D(\pi_{X|W}\|\nu_{X}|\pi_{W})\leq\alpha}$
(3.17) $\qquad\inf_{\pi_{Y|W}:\mathsf{W}_{p}^{p}(\pi_{X|W},\pi_{Y|W}|\pi_{W})\leq\tau}D(\pi_{Y|W}\|\nu_{Y}|\pi_{W}).$
In particular, if further assume $\nu=\nu_{X}=\nu_{Y}$, then (3.16) holds for any $\alpha\geq 0,\tau\geq 0$.
Proof.
We prove this theorem by using Theorem 8. By Statement 2 of Theorem 8, for any $(\alpha,\tau)$ in the interior of $\mathrm{dom}\psi:=\left\{(\alpha,\tau):\psi(\alpha,\tau)<\infty\right\}$, it holds that
[…]

### v4

@@ 3.4. Proof of Proposition 3
To show that the ratio $f(R)=\frac{V_{n}(\mathbf{p},R)}{V_{n}^{(K)}(R)}$ is non-increasing, we utilize the properties of the Bishop-Gromov comparison theorem and the behavior of ratios of integrals.
Step 1: As shown above, the ratio of the areas of geodesic spheres $\frac{A(p,r)}{A^{(K)}(r)}$ is a non-increasing function of $r$. Let $\rho=(\rho_{1},\dots,\rho_{n})$ be a vector in the integration domain. For any scalar $r>0$ and a unit vector $\omega\in S_{+}^{n-1}=\{\omega\in S^{n-1}:\omega_{i}\geq 0\text{\ for\ all\ }i\}$, let $\rho=r\omega$. The product ratio is: $\frac{\prod_{i=1}^{n}A(p_{i},r\omega_{i})}{\prod_{i=1}^{n}A^{(K)}(r\omega_{i})}=\prod_{i=1}^{n}\frac{A(p_{i},r\omega_{i})}{A^{(K)}(r\omega_{i})}$. Since each term $\frac{A(p_{i},r\omega_{i})}{A^{(K)}(r\omega_{i})}$ is non-increasing in $r$ (because $r\omega_{i}$ increases with $r$ for $\omega_{i}\geq 0$), the product of these non-negative non-increasing functions is also non-increasing in $r$.
Step 2: We express the volumes $V_{n}(R)$ and $V_{n}^{(K)}(R)$ using polar coordinates in the $\rho$-space $\mathbb{R}^{n}$$V_{n}(R)=V_{n}(\mathbf{p},R)=\int_{0}^{R}r^{n-1}\left(\int_{S_{+}^{n-1}}\prod A(p_{i},r\omega_{i})d\omega\right)dr$. Let $G(r)=G(\mathbf{p},r)=r^{n-1}\int_{S_{+}^{n-1}}\prod A(p_{i},r\omega_{i})d\omega$ and $H(r)=r^{n-1}\int_{S_{+}^{n-1}}\prod A^{(K)}(r\omega_{i})d\omega$. From Step 1, the ratio of the integrands $\frac{\prod A(p_{i},r\omega_{i})}{\prod A^{(K)}(r\omega_{i})}$ is non-increasing in $r$ for every $\omega$. Consequently, the ratio of the integrals over the angular components, $\frac{G(r)}{H(r)}$, is also non-increasing.
Step 3: A standard lemma in analysis states that if $G(r)$ and $H(r)$ are positive functions such that $\frac{G(r)}{H(r)}$ is non-increasing, then the ratio of their cumulative integrals $F(R)=\frac{\int_{0}^{R}G(r)dr}{\int_{0}^{R}H(r)dr}$ is also non-increasing. To see this, calculate the derivative $F^{\prime}(R)$: $F^{\prime}(R)=\frac{G(R)\int_{0}^{R}H(r)dr-H(R)\int_{0}^{R}G(r)dr}{(\int_{0}^{R}H(r)dr)^{2}}$. Since $\frac{G(r)}{H(r)}\geq\frac{G(R)}{H(R)}$ for all $r\leq R$, we have $H(R)G(r)\geq G(R)H(r)$. Integrating both sides w.r.t $r$ from 0 to $R$ yields $H(R)\int_{0}^{R}G(r)dr\geq G(R)\int_{0}^{R}H(r)dr$. Thus, $F^{\prime}(R)\leq 0$. The function $f(R)=\frac{V_{n}(R)}{V_{n}^{(K)}(R)}$ is non-increasing for $R>0$. ∎
In a model space $M^{(K)}$ of dimension $k=\mathrm{dim}(\mathcal{M})+N$, the area $A^{(K)}(\rho)$ of a geodesic sphere of radius $\rho$ is given by
$A^{(K)}(\rho)=S_{k-1}\cdot\mathrm{sn}_{\kappa}(\rho)^{k-1},$
where $S_{k-1}=\frac{2\pi^{k/2}}{\Gamma(k/2)}$ is the area of a unit (k1)(k-1)-sphere (in Euclidean spaces), and the function $\mathrm{sn}_{\kappa}(\rho)$ is defined as
$\mathrm{sn}_{\kappa}(\rho)=\begin{cases}\frac{1}{\sqrt{\kappa}}\sin(\sqrt{\kappa}\rho),&K>0,\\ \rho,&K=0,\\ \frac{1}{\sqrt{|\kappa|}}\sinh(\sqrt{|\kappa|}\rho),&K<0,\end{cases}$
with $\kappa=\frac{K}{k-1}$.
To prove the desired result, it suffices to only consider the weaker case, $K<0$. Recalling the definition of $N_{n}$ in (3.10) and by Lemma 1,
$N_{n}(\sqrt{n}R,\sqrt{n}r)=\sup_{\mathbf{p}}\frac{V_{n}(\mathbf{p},\sqrt{n}R)}{V_{n}(\mathbf{p},\sqrt{n}r)}\leq\frac{V_{n}^{(K)}(\sqrt{n}R)}{V_{n}^{(K)}(\sqrt{n}r)}.$
By large deviations theory (or Laplace’s Method), one can determine the asymptotics of $V_{n}^{(K)}(\sqrt{n}R)$ as follows
$\lim_{n\rightarrow\infty}\frac{1}{n}\ln V_{n}^{(K)}(\sqrt{n}R)=\sqrt{|K|(k-1)}R.$
Hence, substituting $R\leftarrow\lambda r$, we obtain that
$\lim_{r\to\infty}\limsup_{n\to\infty}\frac{1}{nr^{2}}\ln N_{n}(\sqrt{n}\lambda r,\sqrt{n}r)=\lim_{r\to\infty}\frac{1}{r^{2}}\sqrt{|K|(k-1)}(\lambda-1)r=0.$
3.5. Proof of Theorem 8
We now provide the proof of Theorem 7 with the outline given as follows. By Sanov’s theorem, we may restrict most coordinates to compact sets without altering the exponential order of the set’s size, though the remaining coordinates cannot be constrained. Fortunately, in the case that the condition (1.9) holds, the measure $\nu^{\otimes n}$ exhibits stronger concentration than Gaussian measures, which allows us to cover the majority of the mass of the remaining coordinates at only a negligible enlargement cost (i.e., in terms of distance). In other words, the isoperimetric exponent in this case is determined by the one for compact sets and the latter was already characterized in Theorem 7. On the other hand, in the case that the condition (1.10) holds, under the covering bound assumption (Condition 4), the majority of the mass of the remaining coordinates can be covered by subexponentially many small balls with negligible radius. Thus, there is one small ball that dominates others in the sense that it has almost the same mass (up to a subexponential factor) as the whole mass of the remaining coordinates. Covering this small ball will cover the majority of the mass of the remaining coordinates at only a negligible enlargement cost (i.e., in terms of distance). In this case, the isoperimetric exponent is also determined by the one for compact sets characterized in Theorem 7. We next present the detailed proof.
3.5.1. Step 1: Tailoring the Set
Let $A_{n}$ be a sequence of optimal subsets asymptotically attaining the optimal exponent $\limsup_{n\to\infty}E_{n}(\alpha,\tau)$. Specifically, we let $A_{n}$ be a sequence of subsets such that $\nu^{\otimes n}(A_{n})\geq e^{-n\alpha}$ and $\frac{1}{n}\ln\nu^{\otimes n}(A_{n}^{\sqrt{n\tau}})\geq E_{n}(\alpha,\tau)-\frac{1}{n}.$ When there is no ambiguity, for brevity we denote $A=A_{n}$.
Since $\mathcal{X}$ is Polish, any probability measure on it is tight. So, for any $\eta\in(0,1)$, there is a compact set $F=F_{\eta}\subseteq\mathcal{X}$ such that $\nu(F^{c})\leq\eta$. Let $\mathbf{X}=(X_{1},...,X_{n})\sim\nu^{\otimes n}$ and $S_{i}:=\mathbbm{1}_{F^{c}}(X_{i})$, $i\in[n]$. Then, $\mathbf{S}\sim\mathrm{Bern}(\nu(F^{c}))^{\otimes n}$. By Sanov’s theorem, for any $\epsilon\in(\eta,1)$,
$\mathbb{P}\{|\mathbf{S}|>n\epsilon\}\leq e^{-nD(\epsilon\|\nu(F^{c}))}\leq e^{-nD(\epsilon\|\eta)},$
where $|\mathbf{S}|=\sum_{i=1}^{n}S_{i}$ denotes the number of “1” in $\mathbf{S}$, $D(\epsilon\|\eta)=\epsilon\ln\frac{\epsilon}{\eta}+(1-\epsilon)\ln\frac{1-\epsilon}{1-\eta}$ is the binary relative entropy function, and the second inequality follows since $\eta\mapsto D(\epsilon\|\eta)$ is decreasing for $\eta<\epsilon$. Since $\eta\mapsto D(\epsilon\|\eta)$ goes to infinity as $\eta\downarrow 0$, given any $\epsilon>0$, we can choose $\eta=\eta_{\epsilon}$ small enough so that $D(\epsilon\|\eta)>\alpha$. For example, we can choose $\eta=\epsilon e^{-\epsilon^{-2}}$ and choose $\epsilon$ small enough. By basic probabilistic inequalities, we have that
$\mathbb{P}\{\mathbf{X}\in A,\,|\mathbf{S}|\leq n\epsilon\}$ $\geq\mathbb{P}\{\mathbf{X}\in A\}-\mathbb{P}\{|\mathbf{S}|>n\epsilon\}\geq e^{-n\alpha}-e^{-nD(\epsilon\|\eta)},$
which implies that
(3.13) $\mathbb{P}\{\mathbf{X}\in A,\,|\mathbf{S}|\leq n\epsilon\}\geq e^{-n(\alpha+o(1))}$
with $o(1)$ denoting a term vanishing as $n\to\infty$.
Under the condition $|\mathbf{S}|\leq n\epsilon$, the random vector $\mathbf{S}$ takes ${n\choose\leq n\epsilon}:=\sum_{i=1}^{\left\lfloor n\epsilon\right\rfloor}{n\choose i}$ distinct vector values. By Sanov’s theorem, ${n\choose\leq n\epsilon}\leq e^{nH(\epsilon)}$, where $H(\epsilon)=-\epsilon\ln\epsilon-(1-\epsilon)\ln(1-\epsilon)$ is the binary entropy function. Combining this with (3.13) yields that
(3.14) $\max_{|\mathbf{s}|\leq n\epsilon}\mathbb{P}\{\mathbf{X}\in A,\mathbf{S}=\mathbf{s}\}\geq e^{-n(\alpha+H(\epsilon)+o(1))}.$
Let $\mathbf{s}^{*}$ be the optimal $\mathbf{s}$ attaining the maximum in the above equation. By rearrangement of the coordinates of $\mathbf{s}^{*}$, we assume $\mathbf{s}^{*}=(0,...,0,1,...,1)$ such that $|\mathbf{s}^{*}|\leq n_{1}:=n\epsilon$. (For brevity, $n\epsilon$ is assumed to be an integer, otherwise, denote $n_{1}:=\left\lfloor n\epsilon\right\rfloor$ and our proof still works.) Let $n_{0}=n-n_{1}=(1-\epsilon)n$. Then, (3.14) implies that $A\cap(F^{n_{0}}\times\mathcal{X}^{n_{1}})$ has probability at least $e^{-n(\alpha+H(\epsilon)+o(1))}$.
By Chernoff’s inequality, for $\mathbf{X}_{1}\sim\nu^{\otimes n_{1}}$,
$\mathbb{P}\{d^{2}(\mathbf{X}_{1},\mathbf{v}_{0})>n\beta\}\leq e^{-n_{1}\left(\lambda\beta/\epsilon-\Psi(\lambda)\right)}=e^{-n\left(\lambda\beta-\epsilon\Psi(\lambda)\right)},$
where $\mathbf{v}_{0}=(x_{0},...,x_{0})\in\mathcal{X}^{n_{1}}$, $\Psi(\lambda):=\ln\mathbb{E}_{\nu}[e^{\lambda d^{2}(X,x_{0})}]$ is the logarithmic moment generating function of $d^{2}(X,x_{0})$, and $\lambda$ is chosen such that $\Psi(\lambda)<\infty$ (by Condition 2, such $\lambda$ exists).
In the case that the condition (1.10) holds, we choose $\beta$ large enough so that $\lambda\beta-\epsilon\Psi(\lambda)>\alpha+H(\epsilon)$ (e.g., $\beta=\beta_{\alpha}:=\frac{\alpha+1+\Psi(\lambda)}{\lambda}$). In the case that the condition (1.9) holds, for any small $\beta>0$, we can choose $\lambda$ large enough and $\epsilon$ small enough so that $\lambda\beta-\epsilon\Psi(\lambda)>\alpha+H(\epsilon)$. In both these two cases, $A\cap(F^{n_{0}}\times B_{\sqrt{n\beta}}(\mathbf{v}_{0}))$ has probability at least $e^{-n(\alpha+H(\epsilon)+o(1))}$.
3.5.2. Step 2: Covering Argument
We next quantize the ball $B_{\sqrt{n\beta}}(\mathbf{v}_{0})$ by a covering argument. Specifically, we cover the ball $B_{\sqrt{n\beta}}(\mathbf{v}_{0})$ by using a collection of small balls of radius $\sqrt{n\delta}$ for a given small $\delta>0$. In the case that the condition (1.9) holds, $\beta$ can be arbitrarily small. Thus, in this case, we can set $\beta\leq\delta$ and $B_{\sqrt{n\beta}}(\mathbf{v}_{0})$ is covered by itself.
We next deal with the case that the condition (1.10) holds. In this case, we need the covering bound assumption (Condition 4). By this assumption, for any $\delta>0$, there is a cover of $B_{\sqrt{n\beta}}(\mathbf{v}_{0})$ consisting of balls $B_{\sqrt{n\delta}}(\mathbf{v}_{i})$ with $\mathbf{v}_{i}\in\mathcal{X}^{n_{1}}$ and $1\leq i\leq N$, where $N=e^{n\epsilon^{\prime}}$ with $\epsilon^{\prime}=\epsilon^{\prime}_{n,\epsilon,\beta,\delta}$ satisfying
$\lim_{\epsilon\to 0}\limsup_{n\to\infty}\epsilon^{\prime}_{n,\epsilon,\beta,\delta}=0,\;\forall\beta>\delta>0.$
Thus, there is at least one small ball $B:=B_{\sqrt{n\delta}}(\mathbf{v}_{i^{*}})$ for some $i^{*}$ such that
$\tilde{A}:=A\cap(F^{n_{0}}\times B)$
has probability at least $e^{-n(\alpha+H(\epsilon)+\epsilon^{\prime}+o(1))}$. Note that for fixed $\beta,\delta$, the exponent of this probability approaches $\alpha$, as $n\to\infty$ first and $\epsilon\to 0$ then.
Denote $\tilde{A}_{0}:=\{\mathbf{u}\in F^{n_{0}}:\exists\mathbf{v},(\mathbf{u},\mathbf{v})\in\tilde{A}\}$ as the the projection of $\tilde{A}$ on the first $n_{0}$ components. Then, $\tilde{A}_{0}\times B$ is a good isoperimetric approximation of $A$ in the sense that $\tilde{A}_{0}\times B$ has probability at least $e^{-n(\alpha+H(\epsilon)+\epsilon^{\prime}+o(1))}$ (since $\tilde{A}\subseteq\tilde{A}_{0}\times B$) and
$A^{\sqrt{n\tau}}\supseteq\tilde{A}^{\sqrt{n\tau}}$ $=\bigcup_{\mathbf{x}\in\tilde{A}}B_{\sqrt{n\tau}}(\mathbf{x})$
$\supseteq\bigcup_{(\mathbf{u},\mathbf{v})\in\tilde{A}}B_{\sqrt{n\tau_{0}}}(\mathbf{u})\times B_{\sqrt{n\tau_{1}}+2\sqrt{n\delta}}(\mathbf{v})$
$\supseteq\bigcup_{(\mathbf{u},\mathbf{v})\in\tilde{A}}B_{\sqrt{n\tau_{0}}}(\mathbf{u})\times B_{\sqrt{n\tau_{1}}+\sqrt{n\delta}}(\mathbf{v}_{i^{*}})$
(3.15) $=\tilde{A}_{0}^{\sqrt{n\tau_{0}}}\times B^{\sqrt{n\tau_{1}}},$
where $\tau_{0},\tau_{1}$ are arbitrary nonnegative numbers such that $\tau_{0}+\left(\sqrt{\tau_{1}}+2\sqrt{\delta}\right)^{2}\leq\tau$. Let $\alpha_{0,n}=-\frac{1}{n}\ln\nu^{\otimes n_{0}}(\tilde{A}_{0})$ and $\alpha_{1,n}=-\frac{1}{n}\ln\nu^{\otimes n_{1}}(B)$. By passing to subsequence, we can assume that $\alpha_{0,n}\to\alpha_{0}$ and $\alpha_{1,n}\to\alpha_{1}$ as $n\to\infty$ for some nonnegative numbers $\alpha_{0},\alpha_{1}$ such that $\alpha_{0}+\alpha_{1}\leq\alpha+H(\epsilon)+\epsilon^{\prime}$.
3.5.3. Step 3: Information-Theoretic Bound
We next use information-theoretic techniques to derive bounds. Denote $\nu_{F}=\nu(\cdot|F)$. Applying Theorem 7 with substitution $(n,\alpha,\tau,\nu_{X},\nu_{Y},A)\leftarrow(n_{0},\alpha_{0},\tau_{0},\nu_{F},\nu,\tilde{A}_{0})$ yields that given any $\delta>0$, there is a probability measure $\pi_{XW}\in\mathscr{P}_{\mathrm{c}}(\mathcal{X}\times\{0,1\})$ induced by the set $\tilde{A}_{0}$ such that $D(\pi_{X|W}\|\nu_{F}|\pi_{W})\leq\frac{\alpha_{0}}{1-\epsilon}+\delta$ and for sufficiently large $n$,
(3.16) $\frac{1}{n_{0}}\ln\nu^{\otimes n_{0}}(\tilde{A}_{0}^{\sqrt{n\tau_{0}}})\leq\inf_{\begin{subarray}{c}\pi_{Y|XW}:\mathbb{E}[d^{2}(X,Y)]\leq\tau_{0}-\delta,\\ \mathbb{E}[\mathrm{Var}(d^{2}(X,Y)|X,W)]\leq\kappa\end{subarray}}(1+\delta)D(\pi_{Y|W}\|\nu|\pi_{W})+\delta.$
We choose $\delta$ here as the same $\delta$ in the radius of the small ball described above.
We next derive a multiletter bound for exponent of probability of $B^{\sqrt{n\tau_{1}}}$. Let $\mu_{\mathbf{X}}=\nu^{\otimes n_{1}}(\cdot|B)$ be the conditional probability measure of $\nu^{\otimes n_{1}}$ given $B$. Then,
(3.17) $D(\mu_{\mathbf{X}}\|\nu^{\otimes n_{1}})=-\log\nu^{\otimes n_{1}}(B).$
Let $\mu_{\mathbf{Y}|\mathbf{X}}$ be a transition probability defined from $\mathcal{X}^{n_{1}}$ to $\mathcal{Y}^{n_{1}}$ such that given each $\mathbf{x}$, $\mu_{\mathbf{Y}|\mathbf{X}=\mathbf{x}}$ is concentrated on the cost ball $B_{\sqrt{n\tau_{1}}}(\mathbf{x}):=\{\mathbf{y}:d(\mathbf{x},\mathbf{y})\leq\sqrt{n\tau_{1}}\}$. Then, we have that $\mu_{\mathbf{Y}}:=\mu_{\mathbf{X}}\circ\mu_{\mathbf{Y}|\mathbf{X}}$ is concentrated on $B^{\sqrt{n\tau_{1}}}$, which implies that $\frac{1}{n_{1}}\ln\nu^{\otimes n_{1}}(B^{\sqrt{n\tau_{1}}})\leq\frac{1}{n_{1}}D_{0}(\mu_{\mathbf{Y}}\|\nu^{\otimes n_{1}})\leq\frac{1}{n_{1}}D(\mu_{\mathbf{Y}}\|\nu^{\otimes n_{1}})$. Here $D_{0}(\mu\|\nu):=-\ln\nu\{\frac{\mathrm{d}\mu}{\mathrm{d}\nu}>0\}$ is the Rényi divergence of order 0, which is no greater than the relative entropy $D(\mu\|\nu)$ [71]. Since $\mu_{\mathbf{Y}|\mathbf{X}}$ is arbitrary, we have that
(3.18) $\log\nu^{\otimes n_{1}}(B^{\sqrt{n\tau_{1}}})\leq\inf_{\mu_{\mathbf{Y}|\mathbf{X}}:d(\mathbf{X},\mathbf{Y})\leq t_{1}\;\mu_{\mathbf{XY}}\textrm{-a.e.}}D(\mu_{\mathbf{Y}}\|\nu^{\otimes n_{1}}).$
We now construct a feasible solution to the infimization above. Given $M,\tau_{1},\delta$, and $\mu_{\mathbf{X}}$, let $\mu_{Y|X}$ be a transition probability such that the induced joint distribution $\mu_{\mathbf{X}\mathbf{Y}}=\mu_{\mathbf{X}}\mu_{Y|X}^{\otimes n_{1}}$ satisfies
(3.19) $\theta(\mathbf{x}):=\sum_{i=1}^{n_{1}}\mathbb{E}_{\mu}[d^{2}(x_{i},Y_{i})|x_{i}]\leq n(\tau_{1}-\delta)\;\mu_{\mathbf{X}}\textrm{-a.e.}\textrm{ and }\;\sum_{i=1}^{n_{1}}\mathrm{Var}_{\mu}(d^{2}(x_{i},Y_{i})|x_{i})\leq nM\;\mu_{\mathbf{X}}\textrm{-a.e.}$
In the following, we modify $\mu_{\mathbf{Y}|\mathbf{X}}$ by truncating it into some high-probability set, in order to make it satisfy the constraint $d(\mathbf{X},\mathbf{Y})\leq\sqrt{n\tau_{1}}$ $\mu_{\mathbf{XY}}\textrm{-a.e.}$
Define
$\theta(x)=\mathbb{E}_{\mu}\left[d^{2}(x,Y)|x\right].$
Then, $\theta(\mathbf{x})=\sum_{i=1}^{n_{1}}\theta(x_{i}).$
By Chebyshev’s inequality, it holds that
$\mu\{\mathbf{Y}\notin B_{\sqrt{n\tau_{1}}}(\mathbf{x})|\mathbf{x}\}$ $=\mu\{d(\mathbf{x},\mathbf{Y})>\sqrt{n\tau_{1}}|\mathbf{x}\}\leq\frac{\mathbb{E}_{\mu}\left[(d^{2}(\mathbf{x},\mathbf{Y})-\theta(x_{i}))^{2}|\mathbf{x}\right]}{(n\tau_{1}-\theta(\mathbf{x}))^{2}}$
$=\frac{\sum_{i=1}^{n_{1}}\mathbb{E}_{\mu}\left[(d^{2}(x_{i},Y_{i})-\theta(x_{i}))^{2}|x_{i}\right]}{(n\tau_{1}-\theta(\mathbf{x}))^{2}}$
$=\frac{\sum_{i=1}^{n_{1}}\mathrm{Var}_{\mu}(d^{2}(x_{i},Y_{i})|x_{i})}{(n\tau_{1}-\theta(\mathbf{x}))^{2}}\leq\frac{M}{n\delta^{2}}=:\epsilon_{n}.$
Here given $M,\delta$, the sequence $\epsilon_{n}$ vanishes if we let $n\to\infty$.
Denote $\hat{\mu}_{\mathbf{Y}|\mathbf{X}}$ as a transition probability given by
$\hat{\mu}_{\mathbf{Y}|\mathbf{X}=\mathbf{x}}=\mu_{\mathbf{Y}|\mathbf{X}=\mathbf{x}}(\cdot|B_{\sqrt{n\tau_{1}}}(\mathbf{x})),\;\forall\mathbf{x}.$
[…]